# Large Language Models for Computer-Aided Design: A Survey

### v1

@@ 5.2. CAD Code Generation
Many LLMs, such as Codex (Chen et al., 2021), are pre-trained on vast code corpora, enabling them to generate code with a high degree of accuracy. Thus, a prominent application of LLMs in the CAD field is the generation of CAD code.
In this direction, Li et al. (Li et al., 2024b, 2025) explored the use of multimodal LLMs for 3D CAD generation, leveraging various data formats, including textual descriptions, images, sketches, and ground truth 3D shapes. In this work, LLMs were employed to generate CAD programs from these inputs, which were subsequently parsed into 3D shapes. Two models were evaluated: GPT-4 (Achiam et al., 2023) and GPT-4V (OpenAI, 2023). To improve the quality of the generated CAD code, the authors developed a debugger that iteratively refined the generated code until it successfully executed. The results showed that GPT-4V outperformed GPT-4, especially when using text-only input. It also surpassed other input combinations (e.g., text $+$ sketch, text $+$ image, and text $+$ sketch $+$ image) in terms of parsing rate and intersection over union.
@@
Sun et al. (Sun et al., 2025) recently conducted a fine-tuning experiment on LLMs, comparing them to the baseline GPT-4 (Achiam et al., 2023) without fine-tuning. They used the dataset from (Li et al., 2024b) and implemented four different sampling strategies for fine-tuning data generation. The first strategy randomly selected 60 code templates, the second ensured diversity in code length, the third focused on the shortest lengths, and the fourth on the longest. Overall, the second strategy achieved the highest average parsing rate, while the third performed best in terms of intersection over union. All four fine-tuned LLMs outperformed the baseline GPT-4 in most cases. In (Du et al., 2024a), Du et al. developed BlenderLLM, a script generation model that underwent SFT and iterative self-improvement. They fine-tuned the Qwen2.5-Coder-7B-Instruct model (Hui et al., 2024) using the BlendNet dataset and created a filter to select high-quality data generated by the model. This approach allowed for iterative optimization through cycles of data generation and model training. On similar lines, Alrashedy et al. (Alrashedy et al., 2024) introduced CADCodeVerify, a framework designed to iteratively verify and refine 3D objects generated from CAD code. CADCodeVerify employs a VLM to assess the generated object by answering a set of questions and correcting deviations. They also introduced a benchmark for CAD code generation, containing 200 prompts paired with expert-annotated scripting code. The framework’s feedback loop involved two steps: generating question-answer pairs and producing feedback without human intervention. The study compared the performance of three LLMs: GPT-4 (Achiam et al., 2023), Gemini (Team et al., 2023), and Code LLaMA (Rozière et al., 2023).
@@
In (Badagabettu et al., 2024), Badagabettu et al. presented Query2CAD, a framework that generated CAD designs by using LLMs to produce executable CAD macros. The process involved the user query being fed into a robust LLM, such as GPT-3.5 Turbo or GPT-4 Turbo, which generated a Python macro. In cases where errors occurred, the error messages and Python code were provided to the LLM, which then generated corrected, executable code. The framework also utilized BLIP2 (Li et al., 2023b) to generate captions for isometric views of CAD models, which were then passed onto the LLM to improve code generation through self-refinement loops. Yuan et al. (Yuan et al., 2024a) fine-tuned pre-trained LLMs to create their own LLM, called OpenECAD, which integrated the logical, visual, and coding abilities of VLMs. They rendered Boundary Representation (B-rep) data to generate 3D model images from three different views: default, orthographic, and transparent. For the base language models, they selected relatively small models like OpenELM (Mehta et al., 2024), Gemma (Team et al., 2024b), and Phi-2 (Abdin et al., 2024). To endow OpenECAD with multimodal conversational capabilities, they leveraged training methods and datasets from LLaVA (Liu et al., 2023a) and used GPT-4 (Achiam et al., 2023) to generate multimodal image-language instruction data. The model was trained using the TinyLLaVA (Zhou et al., 2024) framework and fine-tuned with the LoRA (Hu et al., 2022) method to enable the VLM to generate CAD code effectively.
In addition, several studies have explored the potential of LLMs across a variety of design and manufacturing tasks. For instance, Makatura et al. (Makatura et al., 2024, 2023) evaluated the performance of LLMs in diverse design domains, including 2D vector graphic design (using SVG and DXF formats), 3D parametric modeling (using Constructive Solid Geometry (CSG) and B-rep formats), and articulated robotics problems (using the Universal Robot Description Format (URDF) and general graph-based representations). Except for URDF, the team initially used GPT-4 (Achiam et al., 2023) to generate code, which was then converted into the corresponding 3D models in the desired format. In (Nelson et al., 2023), GPT-4 (Achiam et al., 2023) was used to generate CAD code, employing a back-and-forth dialogue process. Errors in the generated code and resulting structure were fed back into GPT-4, allowing it to iteratively refine the code. The study demonstrated GPT-4’s remarkable ability to generate unique solutions for constructing and debugging CAD code.
Mallis et al. (Mallis et al., 2024) utilized GPT-4o (Hurst et al., 2024) to generate CAD code actions expressed in Python. The generated code was executed in CAD software, and the output was concatenated with the input context before being fed back to GPT-4o for the next iterative step. Rukhovich et al. (Rukhovich et al., 2024) proposed CAD-Recode, a solution for CAD reverse engineering. They fine-tuned an LLM to map input point clouds into CAD sketch-extrude sequences represented as Python code. The LLM used was Qwen2-1.5B (Yang et al., 2024b), a relatively small model. CAD-Recode was augmented with a lightweight, trainable point cloud projector. The framework also utilized GPT-4o (Hurst et al., 2024) to refactor the generated code in real time via interactive sliders.
@@
Jones et al. (Jones et al., 2025) worked on generative CAD design with GPT-4o (Hurst et al., 2024) (without fine-tuning) and a solver-aided, domain-specific language: AI design language (AIDL). The process involved prompting GPT-4o with detailed descriptions of AIDL, including its syntax and available geometry types. The LLM then generated a complete AIDL program prompted by manually designed example programs, which was executed to create the desired model. If errors occurred during execution, they were fed back to GPT-4o, which would correct them. Naik (Naik, 2024) demonstrated the use of LLMs for creating extraction queries in SQL for geometry tagging, streamlining the tool’s functionality by allowing users to provide natural language input instead of complex code. Kienle et al. (Kienle et al., 2024) designed QueryCAD, a deep learning-based question-answering system for CAD models. QueryCAD employed a code-writing LLM with in-context prompting to generate executable code for multi-view CAD part segmentation. It was integrated into MetaWizard, a robot program synthesis system (Alt et al., 2024), to automatically generate industrial robot programs based on natural language task specifications. The team primarily used GPT-4o (Hurst et al., 2024) and compared its performance with other LLaMA models (e.g., LLaMA-3 13B, LLaMA-3.1 8B, and LLaMA-3.1 405B) (Grattafiori et al., 2024).
@@
In (Ocker et al., 2025), Ocker et al. developed a VLM-based multi-agent system for CAD model generation. This system utilized multiple LLM-based agents to interpret sketches, images, and textual descriptions, generate CAD models, and verify the designs. The study proposed strategies for overcoming the well-known spatial reasoning limitations of VLMs, using GPT-4o (Hurst et al., 2024) as the core LLM. In another recent work, Deng et al. (Deng et al., 2024) applied GPT-4 (Achiam et al., 2023) to address three sub-tasks: parametric computation, instruction sequence construction, and coding model scripting. The process involved providing GPT-4 with design requirements, performing iterative design computations, and generating the necessary code by parsing a JSON file to formalize natural language instructions into executable commands.
@@
CAD code generation remains one of the most crucial applications of LLMs in CAD. These generated code can take various forms, such as Python code, SQL code, CAD software-specific code, and even Java or C+++++ + code. Users can generate the desired code formats based on their specific domain knowledge. Figure 2 illustrates the typical CAD code generation pipeline employed by most state-of-the-art approaches.
@@ 5.3. Parametric CAD Generation
Parametric CAD generation focuses on creating parametric data for CAD rather than code. In this application, LLMs generate parametric sequences, often formatted as JSON files.
In the direction of parametric CAD generation, Wu et al. (Wu et al., 2024, 2023b) utilized pre-trained LLMs to efficiently manipulate engineering sketches, integrating sketch primitive sequences and images for parametric CAD generation. This included CAD autocompletion, CAD autoconstraint, and image-conditioned generation. They employed CodeT5+ (Wang et al., 2023c) as the text encoder and decoder for sketch primitive sequences. To assess the general LLM’s capability in CAD autocompletion, they fine-tuned ChatGPT on three subsets, demonstrating that ChatGPT outperformed other baselines. Similarly, Li et al. (Li et al., 2024a) explored text-to-3D parametric CAD generative modeling. During training, both parametric CAD sequences and textual descriptions were used as inputs, and the model generated parametric CAD sequences. At the inference stage, only textual descriptions were provided to generate parametric CAD sequences.
@@
In (Xu et al., 2025), Xu et al. proposed CAD-MLLM, a system designed to generate parametric CAD models conditioned on multimodal inputs such as text, images, and point clouds. For images, they rendered multi-view images from eight fixed perspectives. For point clouds, they randomly sampled points at different ratios and recorded their corresponding normal information. They fine-tuned the open-source LLM Vicuna-7B (Chiang et al., 2023) and used LoRA (Hu et al., 2022) during training to minimize learnable parameters. You et al. (You et al., 2024) focused on reverse engineering 3D CAD models from images. They first used the GPT-4V (OpenAI, 2023) foundation model to predict a global discrete base structure, extracting semantic information from images. The model identified semantic parts from the image before generating CAD sequences for each part. They then built a Transformer model to predict continuous attribute values based on the discrete structure with semantics.
Zhang et al. (Zhang et al., 2025) applied LLMs for controllable generation across various CAD construction hierarchies. They initially converted a CAD model into structured text. They then fine-tuned an LLM to develop a unified model for controllable CAD generation. During training, a hierarchy-aware field in the CAD text was masked, and LLMs were tasked with predicting the masked field. During inference, users could specify the part to modify by masking it and inputting it into the LLM for generating new CAD models. Wang et al. (Wang et al., 2025a) proposed a CAD synthesis method using a multimodal LLM with enhanced spatial reasoning capabilities. They input either a single image or text and mapped the 3D space into 1D using a tokenization method to improve spatial reasoning. They employed LLaVA-1.5 7B (Liu et al., 2024c) as their base model, with pre-trained Vicuna (Chiang et al., 2023) as their foundation, which was built on LLaMA-2 (Touvron et al., 2023b). The 3D CAD model was represented in JSON format, capturing key modeling commands and parameters in the order of CAD construction, based on the DeepCAD dataset (Wu et al., 2021).
@@
Utilizing the LLaMA-3 8B Instruct (Grattafiori et al., 2024) LLM as the backbone, Wang et al. (Wang et al., 2025b) introduced a CAD-Fusion framework. It alternated between two training stages: the sequential learning stage, which trained LLMs using ground truth parametric sequences, and the visual feedback stage, which rewarded parametric sequences that rendered visually preferred objects and penalized those that did not. Makatura et al. (Makatura et al., 2024, 2023) created a design space defined by parametric design and the bounds of these parameters, encompassing a range of potential designs. When prompted with lower and upper bounds for parameters, LLMs suggested values based on typical proportions for the designed object. While the absolute scale was arbitrary, the proposed bounds were semantically reasonable and proportionate. In addition to creating design spaces, they also used GPT-4 (Achiam et al., 2023) to generate XML-format data rather than code, focusing on pre-existing designs in formats like URDF.

### v2

@@ 5.2. CAD Code Generation
Many LLMs, such as Codex (Chen et al., 2021), are pre-trained on vast code corpora, enabling them to generate code with a high degree of accuracy. Thus, a prominent application of LLMs in the CAD field is the generation of CAD code. In this direction, Li et al. (Li et al., 2024c, 2025c) explored the use of multimodal LLMs for 3D CAD generation, leveraging various data formats, including textual descriptions, images, sketches, and ground truth 3D shapes. In this work, LLMs were employed to generate CAD programs from these inputs, which were subsequently parsed into 3D shapes. Two models were evaluated: GPT-4 (Achiam et al., 2023) and GPT-4V (OpenAI, 2023b). To improve the quality of the generated CAD code, the authors developed a debugger that iteratively refined the generated code until successful execution was achieved. The results showed that GPT-4V outperformed GPT-4, especially when using text-only input. It also surpassed other input combinations (e.g., text ++ sketch, text ++ image, and text ++ sketch ++ image) in terms of parsing rate and intersection over union. Sun et al. (Sun et al., 2025) recently conducted a fine-tuning experiment on LLMs, comparing them to the baseline GPT-4 (Achiam et al., 2023) without fine-tuning. They used the dataset from (Li et al., 2024c) and implemented four different sampling strategies for fine-tuning data generation. The first strategy randomly selected 60 code templates, the second ensured diversity in code length, the third focused on the shortest lengths, and the fourth on the longest. Overall, the second strategy achieved the highest average parsing rate, while the third performed best in terms of intersection over union. All four fine-tuned LLMs outperformed the baseline GPT-4 in most cases. In (Du et al., 2024), Du et al. developed BlenderLLM, a script generation model that underwent SFT and iterative self-improvement. They fine-tuned the Qwen2.5-Coder-7B-Instruct model (Hui et al., 2024) using the BlendNet dataset and created a filter to select high-quality data generated by the model. This approach allowed for iterative optimization through cycles of data generation and model training. On similar lines, Alrashedy et al. (Alrashedy et al., 2024) introduced CADCodeVerify, a framework designed to iteratively verify and refine 3D objects generated from CAD code. CADCodeVerify employs a VLM to assess the generated object by answering a set of questions and correcting deviations. They also introduced a benchmark for CAD code generation, containing 200 prompts paired with expert-annotated scripting code. The framework’s feedback loop involved two steps: generating question-answer pairs and producing feedback without human intervention. The study compared the performance of three LLMs: GPT-4 (Achiam et al., 2023), Gemini (Team et al., 2023), and Code LLaMA (Rozière et al., 2023).
@@
In (Badagabettu et al., 2024), Badagabettu et al. presented Query2CAD, a framework that generated CAD designs by using LLMs to produce executable CAD macros. The process involved the user query being fed into a robust LLM, such as GPT-3.5 Turbo (OpenAI, 2023a) or GPT-4 Turbo (OpenAI, 2023c), which generated a Python macro. In cases where errors occurred, the error messages and Python code were provided to the LLM, which then generated corrected, executable code. The framework also utilized BLIP2 (Li et al., 2023a) to generate captions for isometric views of CAD models, which were then passed onto the LLM to improve code generation through self-refinement loops. Yuan et al. (Yuan et al., 2024b) fine-tuned pre-trained LLMs to create their own LLM, called OpenECAD, which integrated the logical, visual, and coding abilities of VLMs. They rendered Boundary Representation (B-rep) data to generate 3D model images from three different views: default, orthographic, and transparent. For the base language models, they selected relatively small models like OpenELM (Mehta et al., 2024), Gemma (Team et al., 2024b), and Phi-2 (Abdin et al., 2024). To endow OpenECAD with multimodal conversational capabilities, they leveraged training methods and datasets from LLaVA (Liu et al., 2023a) and used GPT-4 (Achiam et al., 2023) to generate multimodal image-language instruction data. The model was trained using the TinyLLaVA (Zhou et al., 2024) framework and fine-tuned with the LoRA (Hu et al., 2022) method to enable the VLM to generate CAD code effectively. In addition, several studies have explored the potential of LLMs across a variety of design and manufacturing tasks. For instance, Makatura et al. (Makatura et al., 2024, 2023) evaluated the performance of LLMs in diverse design domains, including 2D vector graphic design (using SVG and DXF formats), 3D parametric modeling (using Constructive Solid Geometry (CSG) and B-rep formats), and articulated robotics problems (using the Universal Robot Description Format (URDF) and general graph-based representations). Except for URDF, the team initially used GPT-4 (Achiam et al., 2023) to generate code, which was then converted into the corresponding 3D models in the desired format. In (Nelson et al., 2023), GPT-4 (Achiam et al., 2023) was used to generate CAD code, employing a back-and-forth dialogue process. Errors in the generated code and resulting structure were fed back into GPT-4, allowing it to iteratively refine the code. The study demonstrated GPT-4’s remarkable ability to generate unique solutions for constructing and debugging CAD code.
@@
Mallis et al. (Mallis et al., 2025) utilized GPT-4o (Hurst et al., 2024) to generate CAD code actions expressed in Python. The generated code was executed in CAD software, and the output was concatenated with the input context before being fed back to GPT-4o for the next iterative step. Rukhovich et al. (Rukhovich et al., 2025) proposed CAD-Recode, a solution for CAD reverse engineering. They fine-tuned an LLM to map input point clouds into CAD sketch-extrude sequences represented as Python code. The LLM used was Qwen2-1.5B (Yang et al., 2024a), a relatively small model. CAD-Recode was augmented with a lightweight, trainable point cloud projector. The framework also utilized GPT-4o (Hurst et al., 2024) to refactor the generated code in real time via interactive sliders. Jones et al. (Jones et al., 2025) worked on generative CAD design with GPT-4o (Hurst et al., 2024) (without fine-tuning) and a solver-aided, domain-specific language: AI design language (AIDL). The process involved prompting GPT-4o with detailed descriptions of AIDL, including its syntax and available geometry types. The LLM then generated a complete AIDL program prompted by manually designed example programs, which was executed to create the desired model. If errors occurred during execution, they were fed back to GPT-4o, which would correct them. Naik (Naik, 2024) demonstrated the use of LLMs for creating extraction queries in SQL for geometry tagging, streamlining the tool’s functionality by allowing users to provide natural language input instead of complex code. Four LLMs were compared: Gemini Pro (Team et al., 2023), GPT-3.5 Turbo (OpenAI, 2023a), GPT-4, Claude 3 Sonnet (Anthropic, 2024b). Kienle et al. (Kienle et al., 2025) designed QueryCAD, a deep learning-based question-answering system for CAD models. QueryCAD employed a code-writing LLM with in-context prompting to generate executable code for multi-view CAD part segmentation. It was integrated into MetaWizard, a robot program synthesis system (Alt et al., 2024), to automatically generate industrial robot programs based on natural language task specifications. The team primarily used GPT-4o (Hurst et al., 2024) and compared its performance with other LLaMA models (e.g., LLaMA-3 13B, LLaMA-3.1 8B, and LLaMA-3.1 405B) (Grattafiori et al., 2024).
@@
In (Ocker et al., 2025), Ocker et al. developed a VLM-based multi-agent system for CAD model generation. This system utilized multiple LLM-based agents to interpret sketches, images, and textual descriptions, generate CAD models, and verify the designs. The study proposed strategies for overcoming the well-known spatial reasoning limitations of VLMs, using GPT-4o (Hurst et al., 2024) as the core LLM. In another recent work, Deng et al. (Deng et al., 2024) applied GPT-4 (Achiam et al., 2023) to address three sub-tasks: parametric computation, instruction sequence construction, and coding model scripting. The process involved providing GPT-4 with design requirements, performing iterative design computations, and generating the necessary code by parsing a JSON file to formalize natural language instructions into executable commands. Doris et al. (Doris et al., 2025a) introduced CAD-Coder, a fine-tuned open-source VLM for the generation of CAD code based on LLaVA-1.5 (Liu et al., 2023a). Given image inputs, the model produced readily executable Python CadQuery code. They compared their model against both closed-source and open-source VLMs, including GPT-o1 (Jaech et al., 2024), GPT-4.5 (OpenAI, 2025c), and Gemini 2.0 (Pichai et al., 2024) as closed-source, and InternVL2_5-78B-MPO (Wang et al., 2024c), Ovis2-34B (Lu et al., 2024b), and Qwen2.5-VL-72B (Team, 2025) as open-source. Xie et al. (Xie and Ju, 2025) proposed generating CadQuery code directly from text using LLMs. They fine-tuned six open-source models of varying sizes, including CodeGPT-small (Microsoft, 2020), GPT-2 medium and GPT-2 large (Radford et al., 2019), Gemma3-1B (Team et al., 2025), Qwen2.5-3B (Qwen et al., 2025), and Mistral-7B (Jiang et al., 2023). All models were fine-tuned via full-parameter supervised learning, except Mistral-7B, which adopted parameter-efficient fine-tuning using LoRA (Hu et al., 2022). Kolodiazhnyi et al. (Kolodiazhnyi et al., 2025) proposed Cadrille, a CAD reconstruction model that jointly processed point clouds, images, and text, built on top of Qwen2-VL-2B (Wang et al., 2024b). They used a two-stage pipeline: SFT on large-scale procedurally generated data, followed by RL with online feedback, and showed that online RL outperformed offline alternatives.
Furthermore, in (Lv and Bao, 2025), the authors employed Qwen2.5-Coder-7B to generate CAD programs. Likewise, Rosnitschek et al. (Rosnitschek et al., 2025) employed GPT-2 and CodeGen-350M for CAD code generation. In the same vein, Madireddy et al. (Madireddy et al., 2025) used BIM models and regulatory documents as input data and leveraged the capabilities of LLMs to generate Python scripts tailored for automated rule checking. Eight LLMs were evaluated: ChatGPT 4.0, Claude Sonnet 3.5, LLaMA 3.1, Microsoft Copilot, Gemini, Perplexity.AI, Grok, and DeepSeek. Similarly, in (Paviot et al., 2025), authors extracted B-Rep geometry from a CAD model and then used a dedicated API to generate a more concise and readable program. In addition, Campos et al. (Campos et al., 2025) tackled the text-to-SQL task for a specific type of real-world relational database containing data extracted from engineering CAD files. They introduced a task-specific prompt strategy and evaluated models including GPT-4o, LLaMA 3.1 8B, LLaMA 3.1 Instruct 70B, and Gemma 2 27B. Moreover, Chen et al. (Chen et al., 2025) proposed a multimodal LLM-based framework consisting of a feedback generator and a code editor, which performed geometric operations by editing code for 3D object reconstruction. The framework employed two LLMs: Qwen2-VL (Wang et al., 2024b) and LLaVA-OV (Li et al., 2024a). In (Guan et al., 2025), Guan et al. reformulated the text-to-CAD task as generating CadQuery code from natural language descriptions. Their approach consisted of two stages: supervised fine-tuning for CAD code generation and reinforcement learning with CAD-specific reward functions, using Qwen2.5-7B-Instruct (Qwen et al., 2025) as the base model. Panta et al. (Panta et al., 2025) also generated a Python script by importing the required libraries, which specified the parameters of the CAD model using the GPT-4o LLM.
@@
CAD code generation remains one of the most crucial applications of LLMs in CAD. These generated code can take various forms, such as Python code, SQL code, CAD software-specific code, and even Java or C++++ code. Users can generate the desired code formats based on their specific domain knowledge. Figure 2 illustrates the typical CAD code generation pipeline employed by most state-of-the-art approaches.
@@ 5.3. Parametric CAD Generation
Parametric CAD generation focuses on creating parametric data for CAD rather than code. In this application, LLMs generate parametric sequences, often formatted as JSON files. (Zhou and D. Camba, 2025) provides a comprehensive review of multimodal LLMs for parametric CAD. In the direction of parametric CAD generation, Wu et al. (Wu et al., 2024b, 2023b) utilized pre-trained LLMs to efficiently manipulate engineering sketches, integrating sketch primitive sequences and images for parametric CAD generation. This included CAD autocompletion, CAD autoconstraint, and image-conditioned generation. They employed CodeT5+ (Wang et al., 2023c) as the text encoder and decoder for sketch primitive sequences. To assess the general LLM’s capability in CAD autocompletion, they fine-tuned ChatGPT on three subsets, demonstrating that ChatGPT outperformed other baselines. In (Xu et al., 2025), Xu et al. proposed CAD-MLLM, a system designed to generate parametric CAD models conditioned on multimodal inputs such as text, images, and point clouds. For images, they rendered multi-view images from eight fixed perspectives. For point clouds, they randomly sampled points at different ratios and recorded their corresponding normal information. They fine-tuned the open-source LLM Vicuna-7B (Chiang et al., 2023) and used LoRA (Hu et al., 2022) during training to minimize learnable parameters. You et al. (You et al., 2025) focused on reverse engineering 3D CAD models from images. They first used the GPT-4V (OpenAI, 2023b) foundation model to predict a global discrete base structure, extracting semantic information from images. The model identified semantic parts from the image before generating CAD sequences for each part. They then built a Transformer model to predict continuous attribute values based on the discrete structure with semantics.
@@
Zhang et al. (Zhang et al., 2025b) applied LLMs for controllable generation across various CAD construction hierarchies. They initially converted a CAD model into structured text. They then fine-tuned an LLM, LLaMA-3-8B (Grattafiori et al., 2024), to develop a unified model for controllable CAD generation. During training, a hierarchy-aware field in the CAD text was masked, and LLMs were tasked with predicting the masked field. During inference, users could specify the part to modify by masking it and inputting it into the LLM for generating new CAD models. Wang et al. (Wang et al., 2025b) proposed a CAD synthesis method using a multimodal LLM with enhanced spatial reasoning capabilities. They input either a single image or text and mapped the 3D space into 1D using a tokenization method to improve spatial reasoning. They employed LLaVA-1.5 7B (Liu et al., 2024c) as their base model, with pre-trained Vicuna (Chiang et al., 2023) as their foundation, which was built on LLaMA-2 (Touvron et al., 2023b). The 3D CAD model was represented in JSON format, capturing key modeling commands and parameters in the order of CAD construction, based on the DeepCAD dataset (Wu et al., 2021). Utilizing the LLaMA-3 8B Instruct (Grattafiori et al., 2024) LLM as the backbone, Wang et al. (Wang et al., 2025a) introduced a CAD-Fusion framework. It alternated between two training stages: the sequential learning stage, which trained LLMs using ground truth parametric sequences, and the visual feedback stage, which rewarded parametric sequences that rendered visually preferred objects and penalized those that did not. Makatura et al. (Makatura et al., 2024, 2023) created a design space defined by parametric design and the bounds of these parameters, encompassing a range of potential designs. When prompted with lower and upper bounds for parameters, LLMs suggested values based on typical proportions for the designed object. While the absolute scale was arbitrary, the proposed bounds were semantically reasonable and proportionate. In addition to creating design spaces, they also used GPT-4 (Achiam et al., 2023) to generate XML-format data rather than code, focusing on pre-existing designs in formats like URDF. Li et al. (Li et al., 2025a) proposed CAD-Llama, a comprehensive framework that adapted open-source LLMs for generating CAD sequences. They introduced a code-like format to translate parametric 3D CAD command sequences into structured parametric CAD code. For training, they employed an adaptive pretraining approach combined with instruction tuning for various downstream tasks, which enabled the LLM to acquire CAD modeling capabilities and adapt to diverse applications, selecting LLaMA3-8B (Grattafiori et al., 2024) as the foundational LLM. Yuan et al. (Yuan et al., 2025) introduced a text-based CAD editing framework, where the input consisted of an editing instruction and the sequence representation of the original CAD model, and the output was the sequence of the edited model. This approach enabled precise modifications through natural language to better reflect real-world user intent. They fine-tuned LLaMA-3-8B-Instruct (Grattafiori et al., 2024) using LoRA.
Likewise, in (Ghosh et al., 2025), Gosh et al. represented customer requirements as text- or sketch-based descriptions of the product to be manufactured and delivered by the manufacturer. Furthermore, Yin et al. (Yin et al., 2025) introduced ShapeGPT, an LLM trained to produce shape token sequences from text, images, or multimodal inputs. The model also supports downstream tasks including shape editing, shape reasoning, and shape completion. Similarly, the authors of (Daareyni et al., 2025) proposed a framework for LLM-assisted CAD using advanced AI systems, specifically the GPT-4o model, enabling users to generate and manipulate 3D models via text commands, visual inputs (e.g., blueprints, images), voice, and other types of general or professional information. Moreover, Rietschel et al. (Rietschel and Steinfeld, 2025) outlined three principles for developing LLM-driven tools in design research and demonstrated these principles through a system capable of generating parametric design scripts via reasoning-enhanced models. In parallel, in (Zhang et al., 2025a), GeoCAD took three inputs: an original CAD model represented in a hierarchical textual format, a selected local part, and user-provided geometric instructions, and generated a new CAD model that modified only the designated local part while closely adhering to the given instructions. In the same vein, Govindarajan et al. (Govindarajan et al., 2026) fine-tuned Qwen 2.5-Coder-14B on their dataset to generate JSON-formatted CAD sequences from natural language prompts, effectively leveraging the pre-trained model’s capabilities without introducing any specialized embedding layers. In (Li et al., 2025d), Li et al. also leveraged the locally deployed open-source inference LLM DeepSeek-R1 for parametric CAD model generation. The initially generated parametric model was rendered into a sequence of step-wise perspective images, which were then processed by a VLM together with the corresponding CoT traces produced by DeepSeek-R1 to assess and subsequently refine the CAD model. Analogously, in (McCarthy et al., 2025), the authors proposed a novel task for studying and benchmarking multimodal refinement in a CAD design setting. One player, the designer, had a target design which they demonstrated via a combination of text and drawing instructions. The other player, the maker, used a programmatic CAD language to carry out the designer’s instructions and attempted to recreate the target design. They had several LLMs, including GPT-4o, GPT-4o-mini, Claude-3.7-Sonnet, and Qwen2.5-VL-7B-Instruct (Bai et al., 2025), assume the role of the human maker. In a similar manner, Panta et al. (Panta et al., 2025) interpreted and analyzed the user-provided specifications and generated a modeling plan outlining the relevant methods and functions of the open-source Python library.